# Knowledge Accumulation in Continually Learned Representations and the Issue of Feature Forgetting

### v3

@@ 4.2 Exclusion baseline
For the first tasks in the training sequence, there is a significant difference between the relative forgetting at the output and the representation, see Figure 3. After training a few new tasks, the output has $\text{rF}\approx 1$, while the representation has a significantly lower rF. In this subsection, we show that this difference can be explained by knowledge accumulation from other tasks. When learning one task improves the representation for another, those tasks have something in common. Even if a model forgets everything it had learned, the common information can be re-learned from the later tasks. After training on all tasks, the final model would still perform better than a model with random weights, which results in rF scores lower than one. To further unravel how knowledge accumulation influences forgetting, we train a new model $f^{-j}$ on the same task sequence as in the previous section, but without task $T_{j}$, see Figure 3(a). Model $f^{-j}$ cannot forget specific information on task $T_{j}$ as it was never trained on it, but it can accumulate common knowledge from other tasks. We define $\text{EXC}_{i,j}$ as the difference between both in models in LP accuracy for task $T_{j}$ after training task $T_{i}$:
@@
Refer to caption (a) Refer to caption (b)
Table 1: Quantitative results of Figure 3(b). The final difference ($\text{EXC}_{19,j}$) and maximal difference ($\text{EXC}_{j,j}$) between model $f$ and its exclusion counterpart $f^{-j}$. (Mean $\pm$ SE, 5 runs) $T_{j}$ $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{7}$ $T_{8}$ $T_{9}$ $\text{EXC}_{19,j}$ 0.001 $\pm$ 0.004 0.002 $\pm$ 0.011 0.008 $\pm$ 0.004 0.018 $\pm$ 0.013 0.023 $\pm$ 0.012 0.010 $\pm$ 0.010 0.012 $\pm$ 0.015 0.005 $\pm$ 0.007 0.012 $\pm$ 0.012 $\text{EXC}_{j,j}$ 0.234 $\pm$ 0.024 0.159 $\pm$ 0.015 0.151 $\pm$ 0.025 0.165 $\pm$ 0.015 0.119 $\pm$ 0.010 0.132 $\pm$ 0.012 0.140 $\pm$ 0.014 0.130 $\pm$ 0.021 0.153 $\pm$ 0.008
@@ 4.3 Feature forgetting reduces knowledge accumulation
Previous sections showed that there is severe feature forgetting in neural networks and that what is not forgotten can also be learned from other tasks. This leads us to the second question of the introduction: what are the consequences of feature forgetting? Is the forgotten information important for that specific task only, or can it also contribute to more general knowledge accumulation? Phrased differently, how much knowledge accumulation would there be if there was no forgetting? To answer this, we consider a baseline that learns continually but has no forgetting, inspired by Vogelstein et al. (2020) and Yan et al. (2021). This ensemble baseline stores a model copy after every task. During evaluation, their representations are concatenated and a linear probe is trained on top of this concatenation. This ensemble learns exactly the same way as a continually finetuned model, but it cannot forget, as the task’s original representation remains intact and can always be recovered by a head. See Figure 4(a) for a schematic of the approach, and Figure 4(b) for the results. Finetuning and the ensemble are exactly the same after training the first task. Knowledge accumulation from the first task aside, our ensemble baseline accumulates almost three times as much knowledge, signified by a 27.4% improvement on a downstream task versus only 9.3% for finetuning. It is interesting that finetuning also accumulates some knowledge during later tasks, albeit far inferior to what is possible when forgetting nothing.
@@
One task improving the performance of another, an effect of knowledge accumulation, is often named as a desideratum of continual learning (Chen & Liu, 2018), yet it persists as a difficult problem. Better representations should also lead to better continual learners. Intuitively, a better representation can make learning easier, reducing the risk of forgetting. With few samples (e.g. a replay memory), a strong representation can also be used to quickly recover past information, sometimes referred to as ‘fast remembering’ (Hadsell et al., 2020; Davari et al., 2022). However, in Appendix D we re-emphasize that even a strong pretrained representation does not always imply mitigation of forgetting. An important step towards accomplishing these goals is preventing feature forgetting as much as possible without preventing models to learn.
@@ 5 Can feature forgetting be prevented?
Over the last years, many methods to alleviate forgetting have been proposed. In this section, we review examples of some of the main families of methods and evaluate how they deal with feature forgetting and knowledge accumulation. The choice of algorithms is not driven by finding the best possible method, but we try to cover the most central ideas, in their simplest form. We test replay with a simple experience replay algorithm with 20 samples per class (ER), parameter regularization using Memory Aware Synapses (MAS) (Aljundi et al., 2018), functional regularization with Learning without Forgetting (LwF) (Li & Hoiem, 2017) and PackNet (Mallya & Lazebnik, 2018) as an architectural method.
Figure 6 shows the relative forgetting of the tested methods and Figure 7 the results of the exclusion baseline for task $T_{6}$. Additionally, Table 2 reports the average learning accuracy (the accuracy on a task immediately after training on that task) and the knowledge accumulation on a downstream task. For all methods, relative forgetting in the representation is at least as bad as at the output level, except for the first task. Relative forgetting in the representation for ER and MAS is lower than for finetuning, yet their knowledge accumulation is similar. One explanation for this is their lower learning accuracy; they learn less in the first place. LwF has even lower relative forgetting, and in contrast to ER and MAS, the exclusion baseline shows that it remembers at least some information specific to the trained task. Its knowledge accumulation is between the ensemble and finetuning, indicating that more task-specific information is retained than with the other methods, but not everything. PackNet is an entirely different story. It does not forget, but what it learns is a lot less useful for other tasks. This is likely because by having masks specific to a single task, the representation for one task is nothing like the representation for another, hindering effective learning and knowledge accumulation.
@@ 6 Discussion and future work
It is possible to think of two tasks that have nothing in common and thus have no knowledge accumulation. Yet, tasks composed of natural images, like those in the benchmarks we used, do share information. An open question is how much of the shared information is at the image level (e.g. detecting edges and shapes) and how much at the semantic level (e.g. being able to recognize a specific feature, like eyes or car tyres). Further unraveling these two might lead to important insights in how neural networks operate. Likewise, in this paper we always used $n$-way classification in both training and testing, with $n$ constant throughout each experiment. The influence of changing problems during training or evaluation (e.g. a different amount of classes per task during evaluation than during training) on the representation quality is something that should be considered in future work.
7 Related work
@@
A number of recent works use linear probing (Alain & Bengio, 2016) to analyze continually learned representations at the penultimate layer (Hu et al., 2022; Davari et al., 2022; Zhang et al., 2022; Chen et al., 2023; Kim & Han, 2023). Considering the amount of forgetting in the continually learning representations, Davari et al. (2022) conclude that this forgetting is less catastrophic than at the output and suggest that no task-critical information is lost. Hu et al. (2022) and Zhang et al. (2022) evaluate the representation’s quality with respect to downstream tasks and arrive at similar findings as Davari et al.. Yoon et al. (2023) consider a continual pre-training setting that allows fine-tuning of all parameters in the model during probing. They report incrementally improving performance on a downstream task, and ascribe it to an increasing transferability of the features learned continually. In contrast to these studies just listed, Kim & Han (2023) show severe forgetting in the representation, but start from models pre-trained on half the respective dataset rather than the more even split considered in Davari et al. (2022). Chen et al. (2023) show that continual learners that have less forgetting at the representation level are better few-shot learners, hinting at the same relation we show with our ensemble baseline.
@@ 8 Conclusion
In this work we studied how deep neural networks learn and forget representations when continually trained on a sequence of image classification tasks. If forgetting is expressed as the proportion of newly gained performance that is forgotten, representations forget about as much as at the output level, except for the first tasks. During the first tasks, information that is shared across tasks is learned, which is not entirely forgotten. The information that is forgotten at the representation level reduces how much knowledge a model accumulates, as exemplified by the ensemble baseline. Finally, we compared the feature forgetting and knowledge accumulation of different types of continual learning methods, whereby we found that functional regularization can prevent a large portion of representation forgetting. We hope that with the work we present here, future continual learning solutions will be evaluated not only on their output performance, but also on their representation quality and how they prevent feature forgetting.
@@ Appendix B Ensemble: further details
To mitigate the influence of the higher dimension, for which it might be easier to find linearly separable features, we add a dimension reduction to lower the dimension back to $k$. We do this by projecting the features of the ensemble on the top-$k$ most significant PCA dimensions. The results are shown in Figure 8. The reduced ensemble performs a bit worse than the full ensemble, yet still significantly better than finetuning.
@@ Appendix D Pre-trained representations
Similarly to representations being attributed an innate robustness to forgetting when learning continually, pre-trained backbones receive this reputation as well (Lee et al., 2023; Zhang et al., 2023). In light of our study, here we also consider the extent to which representations considered “high quality”, such as obtained from large-scale pre-training (Kornblith et al., 2019), impact forgetting at representation level. For this experiment we simply exchange the random initialization that we have been using through the main body with weights from a backbone pre-trained on ImageNet1k  (Deng et al., 2009) 222weights are publicly available in the PyTorch-Torchvision library. An innate robustness to forgetting of the representation should manifest as overall improved performance due to the increased generality and/or quality of the pre-trained features. Note that we exemplarily investigate an extreme case where pre-training (ImageNet1k) coveres an almost identical data distribution as the continual training sequence (Split Mini-Imagenet).
@@
Appendix E Relative forgetting: extra results
@@
Figure 14: Representation and observed forgetting using linear probes for all tasks in Mini-ImageNet using finetuning (except first and last, for which we cannot calculate relative forgetting)
Appendix F Results on CIFAR-100
@@
Figure 16: Representation and observed forgetting using linear probes for all tasks in CIFAR-100 using finetuning (except first and last task, for which we cannot calculate relative forgetting)
@@
Appendix G Evaluation with kNN
In this section we report the most important results of the main paper using kNN instead of using linear probes. This has the benefit that it there are no hyper-parameters to tune and does not depend on the optimization used. We report it here for completeness, and keep the linear probes in the main paper as this is how previous papers reported their results Davari et al. (2022); Cha et al. (2022); Zhang et al. (2022). In general, the results in Figure 20, 21, 23 and 23 follow the same trends as observed in the main paper, with the main difference that the absolute values are lower, likely due to the suboptimalitiy of kNN compared to linear probe optimization.
@@ Appendix G Evaluation with kNN
Figure 21: Representation and observed forgetting using kNN for all tasks in Mini-ImageNet (except last and first, for which we cannot calculate relative forgetting)
Figure 22: Finetune, exclusion, single task and multi task using kNN.
@@
Appendix H Detailed task sequence information

### v4

@@ 4.2 Exclusion baseline
The previous section showed that nearly all performance gained during a task tends to be quickly lost both for the representation and at the output, except for the first tasks in the training sequence. For these tasks, there is a significant difference between the relative forgetting at the output and the representation, see Figure 3. After training a few new tasks, the output has $\text{rF}\approx 1$, while the representation has a significantly lower rF. In this subsection, we show that this difference can be explained by knowledge accumulation from other tasks. When learning one task improves the representation for another, those tasks have something in common. Even if a model forgets everything it had learned, the common information can be re-learned from the later tasks. After training on all tasks, the final model would still perform better than a model with random weights, which results in rF scores lower than one. To further unravel how knowledge accumulation influences forgetting, we train a new model $f^{-j}$ on the same task sequence as in the previous section, but without task $T_{j}$, see Figure 4(a). Model $f^{-j}$ cannot forget specific information on task $T_{j}$ as it was never trained on it, but it can accumulate common knowledge from other tasks. We define $\text{EXC}_{i,j}$ as the difference between both in models in LP accuracy for task $T_{j}$ after training task $T_{i}$:
@@
Table 1: Quantitative results of Figure 4(b). The final difference ($\text{EXC}_{19,j}$) and maximal difference ($\text{EXC}_{j,j}$) between model $f$ and its exclusion counterpart $f^{-j}$. (Mean $\pm$ SE, 5 runs)
$T_{j}$ $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{7}$ $T_{8}$ $T_{9}$
$\text{EXC}_{19,j}$ 0.001 $\pm$ 0.004 0.002 $\pm$ 0.011 0.008 $\pm$ 0.004 0.018 $\pm$ 0.013 0.023 $\pm$ 0.012 0.010 $\pm$ 0.010 0.012 $\pm$ 0.015 0.005 $\pm$ 0.007 0.012 $\pm$ 0.012
$\text{EXC}_{j,j}$ 0.234 $\pm$ 0.024 0.159 $\pm$ 0.015 0.151 $\pm$ 0.025 0.165 $\pm$ 0.015 0.119 $\pm$ 0.010 0.132 $\pm$ 0.012 0.140 $\pm$ 0.014 0.130 $\pm$ 0.021 0.153 $\pm$ 0.008
@@ 4.3 Feature forgetting reduces knowledge accumulation
Previous sections showed that feature forgetting severely impacts newly learned features in neural networks and that what is not forgotten can also be learned from other tasks. This leads us to the second question of the introduction: what are the consequences of feature forgetting? Is the forgotten information important for that specific task only, or can it also contribute to more general knowledge accumulation? Phrased differently, how much knowledge accumulation would there be if there was no forgetting? To answer this, we consider a baseline that learns continually but has no forgetting, inspired by Vogelstein et al. (2020) and Yan et al. (2021). This ensemble baseline stores a model copy after every task. During evaluation, their representations are concatenated and a linear probe is trained on top of this concatenation. This ensemble learns exactly the same way as a continually finetuned model, but it cannot forget, as the task’s original representation remains intact and can always be recovered by a head. See Figure 5(a) for a schematic of the approach, and Figure 5(b) for the results. Finetuning and the ensemble are exactly the same after training the first task. Knowledge accumulation from the first task aside, our ensemble baseline accumulates almost three times as much knowledge, signified by a 27.4% improvement on a downstream task versus only 9.3% for finetuning. It is interesting that finetuning also accumulates some knowledge during later tasks, albeit far inferior to what is possible when forgetting nothing.
@@
One task improving the performance of another, an effect of knowledge accumulation, is often named as a desideratum of continual learning (Chen & Liu, 2018), yet it persists as a difficult problem. Better representations should also lead to better continual learners. If one representation is better than another for data of a new task, it likely needs to change less to reach the same performance after training, thereby reducing the risk of forgetting. With few samples (e.g. a replay memory), a strong representation can also be used to quickly recover past information, sometimes referred to as ‘fast remembering’ (Hadsell et al., 2020; Davari et al., 2022). However, in Appendix D we re-emphasize that even a strong pretrained representation does not always imply mitigation of forgetting. An important step towards accomplishing these goals is preventing feature forgetting as much as possible without preventing models to learn.
@@ 5 Can feature forgetting be prevented?
Over the last years, many methods to alleviate forgetting have been proposed. In this section, we review examples of some of the main families of methods and evaluate how they deal with feature forgetting and knowledge accumulation. The choice of algorithms is not driven by finding the best possible method, but we try to cover the most central ideas, in their simplest form. We test replay with a simple experience replay algorithm with 20 samples per class (ER), parameter regularization using Memory Aware Synapses (MAS) (Aljundi et al., 2018), functional regularization with Learning without Forgetting (LwF) (Li & Hoiem, 2017) and PackNet (Mallya & Lazebnik, 2018) as an architectural method. More variations on the standard replay algorithm are reported in the Appendix.
Figure 6 shows the relative forgetting of the tested methods and Figure 7 the results of the exclusion baseline for task $T_{6}$. Additionally, Table 2 reports the average learning accuracy (the accuracy on a task immediately after training on that task) and the knowledge accumulation on a downstream task. For all methods, relative forgetting in the representation is at least as bad as at the output level, except for the first task. Relative forgetting in the representation for ER and MAS is lower than for finetuning, yet their knowledge accumulation is similar. Adding more samples to the memory in ER does improve knowledge accumulation and reduces overall forgetting, see Appendix. One explanation for this is their lower learning accuracy; they learn less in the first place. LwF has even lower relative forgetting, and in contrast to ER and MAS, the exclusion baseline shows that it remembers at least some information specific to the trained task. Its knowledge accumulation is between the ensemble and finetuning, indicating that more task-specific information is retained than with the other methods, but not everything. PackNet is an entirely different story. It does not forget, but what it learns is a lot less useful for other tasks. This is likely because by having masks specific to a single task, the representation for one task is nothing like the representation for another, hindering effective learning and knowledge accumulation.
@@ 6 Discussion and future work
It is possible to think of two tasks that have nothing in common and thus learning one task does not improve the representation for the other. Yet, tasks composed of natural images, like those in the benchmarks we used, do share information. An open question is how much of the shared information is at the image level (e.g. detecting edges and shapes) and how much at the semantic level (e.g. being able to recognize a specific feature, like eyes or car tyres). Further unraveling these two might lead to important insights in how neural networks operate. Likewise, in this paper we always used $n$-way classification in both training and testing, with $n$ constant throughout each experiment. The influence of changing problems during training or evaluation (e.g. a different amount of classes per task during evaluation than during training) on the representation quality is something that should be considered in future work.
7 Background
In this section we provide more background information related to the results in this paper. They are not necessary to understand our paper, but they may be good starting points to explore the subject further.
@@ 7 Background
A number of recent works use linear probing (Alain & Bengio, 2016) to analyze continually learned representations at the penultimate layer (Hu et al., 2022; Davari et al., 2022; Zhang et al., 2022; Chen et al., 2023; Kim & Han, 2023; Anthes et al., 2023). Considering the amount of forgetting in the continually learning representations, Davari et al. (2022) conclude that this forgetting is less catastrophic than at the output and suggest that no task-critical information is lost. Hu et al. (2022) and Zhang et al. (2022) evaluate the representation’s quality with respect to downstream tasks and arrive at similar findings as Davari et al.. Yoon et al. (2023) consider a continual pre-training setting that allows fine-tuning of all parameters in the model during probing. They report incrementally improving performance on a downstream task, and ascribe it to an increasing transferability of the features learned continually. In contrast to these studies just listed, Kim & Han (2023) show severe forgetting in the representation, but start from models pre-trained on half the respective dataset rather than the more even split considered in Davari et al. (2022). Chen et al. (2023) show that continual learners that have less forgetting at the representation level are better few-shot learners, hinting at the same relation we show with our ensemble baseline.
@@ 8 Conclusion
In this work we studied how deep neural networks learn and forget representations when continually trained on a sequence of image classification tasks. If forgetting is expressed as the proportion of newly gained performance that is forgotten, representations forget about as much as at the output level, except for the first tasks. During the first task, information that is shared across tasks is learned for the first time. Because this shared information is repeated to the model in each task, it is not forgotten. The information that is forgotten at the representation level reduces how much knowledge a model accumulates, as exemplified by the ensemble baseline. Finally, we compared the feature forgetting and knowledge accumulation of different types of continual learning methods, whereby we found that functional regularization can prevent a large portion of representation forgetting. We hope that the insights provided by our work inspire the development of continual learning methods with less feature forgetting and more knowledge accumulation.
@@ Acknowledgments
    Relative forgetting: more results with replay
  • F
@@ Appendix B Ensemble: further details
To mitigate the influence of the higher dimension, for which it might be easier to find linearly separable features, we add a dimension reduction to lower the dimension back to $k$. We do this by projecting the features of the ensemble on the top-$k$ most significant PCA dimensions. The results are shown in Figure 8. The reduced ensemble performs slightly worse than the full ensemble, yet still significantly better than finetuning.
@@ Appendix D Pre-trained representations
Similarly to representations being attributed an innate robustness to forgetting when learning continually, pre-trained backbones receive this reputation as well (Lee et al., 2023; Zhang et al., 2023). In light of our study, here we also consider the extent to which representations considered “high quality”, such as obtained from large-scale pre-training (Kornblith et al., 2019), impact forgetting at representation level. For this experiment we simply exchange the random initialization that we have been using through the main body with weights from a backbone pre-trained on ImageNet1k (Deng et al., 2009). [^2] Weights are publicly available in the PyTorch-Torchvision library. An innate robustness to forgetting of the representation should manifest as overall improved performance due to the increased generality and/or quality of the pre-trained features. Note that we exemplarily investigate an extreme case where pre-training (ImageNet1k) coveres an almost identical data distribution as the continual training sequence (Split Mini-Imagenet).
@@
Appendix E Relative forgetting: more replay results
In Figure 14 we show relative forgetting on the Mini-ImageNet sequence when storing 20, 50 or 100 samples per class, and when adding distillation using the output logits (DER (Buzzega et al., 2020)). Figure 15 shows the exclusion results of $T_{6}$ and Table 3 the amount of knowledge accumulation. Having more samples in the memory reduces both forgetting in the representation and the head and substantially improves knowledge accumulation. Interestingly, the final exclusion difference remains small. This supports our theory. When adding more replay samples the model can learn more from the combination of replay and new task data, but what cannot be learned from this combination of data is still lost. DER is mostly helpful to further reduce forgetting in the head. When storing the output logits for each sample, more information is stored, and it is possible that this reduces the risk of overfitting of the head on the few samples that are in the memory.
(a) Replay – ER (20 samples / class)
(b) Replay – DER Buzzega et al. (20 samples / class)
(c) Replay – ER (50 samples / class)
(d) Replay – ER (100 samples / class)
Figure 14: Relative Forgetting at the output level and the level of representation as in Eq. 1, for $T_{1}$ to $T_{5}$ of more replay variants on the Mini-ImageNet sequence. The influence of storing more samples and adding distillation with soft logits (DER (Buzzega et al., 2020)) is evaluated. (Mean $\pm$ SE, 5 runs)
Figure 15: Task exclusion results of more replay variants on the Mini-ImageNet sequence. The influence of storing more samples and adding distillation with soft logits (DER (Buzzega et al., 2020)) is evaluated. (Mean $\pm$ SE, 5 runs)
Table 3: Learning accuracy averaged over all tasks ($\overline{\text{LP}_{i,i}}$) and knowledge accumulation on a downstream task ($\text{LP}_{19,d}-\text{LP}_{0,d}$), reported when adding more replay samples and with DER (Buzzega et al., 2020). (Mean $\pm$ SE, 5 runs)
Finetune Ensemble ER $(n=20)$ ER $(n=50)$ ER $(n=100)$ DER $(n=20)$
$\overline{\text{LP}_{i,i}}$ 0.860 $\pm$ 0.002 0.864 $\pm$ 0.002 0.815 $\pm$ 0.002 0.814 $\pm$ 0.003 0.821 $\pm$ 0.003 0.826 $\pm$ 0.003
$\text{LP}_{19,d}-\text{LP}_{0,d}$ 0.249 $\pm$ 0.023 0.365 $\pm$ 0.020 0.247 $\pm$ 0.018 0.298 $\pm$ 0.014 0.309 $\pm$ 0.015 0.271 $\pm$ 0.013
Appendix F Relative forgetting: extra results
@@
Figure 16: Representation and observed forgetting using linear probes for all tasks in Mini-ImageNet using finetuning (except first and last, for which we cannot calculate relative forgetting).
Appendix G Results on CIFAR-100
@@
Figure 18: Representation and observed forgetting using linear probes for all tasks in CIFAR-100 using finetuning (except first and last task, for which we cannot calculate relative forgetting).
@@
Appendix H Evaluation with kNN
In this section we report the most important results of the main paper using kNN instead of using linear probes. This has the benefit that it there are no hyper-parameters to tune and does not depend on the optimization used. We report it here for completeness, and keep the linear probes in the main paper as this is how previous papers reported their results Davari et al. (2022); Cha et al. (2022); Zhang et al. (2022). In general, the results in Figure 22, 23, 25 and 25 follow the same trends as observed in the main paper, with the main difference that the absolute values are lower, likely due to the suboptimalitiy of kNN compared to linear probe optimization.
@@ Appendix H Evaluation with kNN
Figure 23: Representation and observed forgetting using kNN for all tasks in Mini-ImageNet (except last and first, for which we cannot calculate relative forgetting).
@@
Appendix I Detailed task sequence information